OpenStudy (anonymous):

6x + 6y + 3z=494 3x + 3y + 3z= 250 in the system of equations above, what is the value of 9x+9y

OpenStudy (anonymous):

subtract the equations from each other\[\huge {(1)-(2)}\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

3x+3y=494-250 then multiply by 3

OpenStudy (anonymous):

why 3

OpenStudy (anonymous):

so that we have 3*3=9 9x+9y=3(494-250)

OpenStudy (anonymous):

k